# The Geometric Reasoner: Manifold-Informed Latent Foresight Search for Long-Context Reasoning

### v1

@@ 3.1 Latent Space State Representation
Given a query $Q$, TGR generates output as a sequence of $L$ chunks $\tau_{1:L}=(\tau_{1},\ldots,\tau_{L})$, each of maximum length $S$. Let $c_{t-1}$ denote the token context preceding chunk $t$. To bound memory during long-horizon decoding, we employ chunk-wise KV cache resets: at each boundary, we rebuild the prefix cache from $c_{t-1}$, discarding earlier history while transmitting long-range state via latent anchors (Appendix LABEL:app:impl). Concretely, $c_{t-1}$ comprises the query and recent token suffix, left-truncated to maximum length $S$.
@@ 3.3 Manifold-Informed Regularization
This penalizes minimal-drift updates only when similarity exceeds $\delta$ (default $\delta=0.2$, see Appendix LABEL:app:hyperparams).
@@ 3.4 Overall Process
This creates a lightweight control interface without parameter updates. We inject at every layer using identical interfaces for candidate evaluation and generation, with fixed random matrices of rank $r$ (default $r=8$) to maintain the training-free setting (Appendix LABEL:app:impl).
@@ 3.5 Vanishing Acceptance of Hard Geometric Constraints
Crucially, $\alpha$ collapses rapidly in high-dimensional latent spaces due to concentration of measure; feasible regions often occupy an exponentially small fraction of the local neighborhood as dimensionality increases (Betancourt, 2017; Goyal and Shetty, 2019; Yallup et al., 2025). Table LABEL:tab:hard_acceptance illustrates that hard filtering becomes computationally prohibitive as dimension increases.
@@ 4.1 Experimental Setup
We use Qwen3 (Yang et al., 2025) as the primary testbed and report main results for the 8B variant; additional scales, 1.7B, 4B and 14B, are deferred to Appendix LABEL:app:additional_results. To study the interaction with training stages, we also evaluate Phi-4-reasoning (SFT Pre-Training) and Phi-4-reasoning-plus (with RL Post-Training) (Abdin et al., 2025).
@@
We report Pass@$k$ for $k\in\mathcal{K}$ and the normalized area under the Pass@$k$ curve, namely AUC as robustness metrics, along with Avg. Tokens, defined as the average end-to-end tokens per problem required to produce $k$ final trajectories (Appendix LABEL:app:exp_grid).
@@
where $\mathcal{K}=\{k_{0}<\dots<k_{M}\}$. For code benchmarks, Pass@$k$ is computed using the unbiased estimator of Chen (2021). For math benchmarks, Pass@$k$ is the empirical success rate that at least one of $k$ samples is correct. We use temperature $0.6$ and the default TGR sampling with $K=8$ candidate anchors per chunk boundary, scored by rapid rollouts ($s=32$ on math tasks and $s=64$ on code tasks); full hyperparameters are listed in Appendix LABEL:app:hyperparams. We set an upper bound of $L=24$ chunks with maximum chunk length $S=512$, while actual generation terminates early by standard stopping criteria; all reported costs are based on realized token consumption.
@@ 4.4 Interpreting TGR Representations
We find that robustness exhibits a consistent geometric signature, with AUC peaking at an intermediate mean geodesic curvature $\kappa$ that is neither overly rigid nor overly erratic and increasing with the effective candidate-set size proxy $N_{\mathrm{eff}}$. This pattern supports the claim that TGR’s advantage stems from maintaining genuinely distinct search paths rather than redundant variations, as illustrated in Figure LABEL:fig:fa_curv_neff).

### v2

@@ 3.1 Latent Space State Representation
Given a query $Q$, TGR generates output as a sequence of $L$ chunks $\tau_{1:L}=(\tau_{1},\ldots,\tau_{L})$, each of maximum length $S$. Let $c_{t-1}$ denote the token context preceding chunk $t$. To bound memory during long-horizon decoding, we employ chunk-wise KV cache resets: at each boundary, we rebuild the prefix cache from $c_{t-1}$, discarding earlier history while transmitting long-range state via latent anchors. Concretely, $c_{t-1}$ comprises the query and recent token suffix, left-truncated to maximum length $S$.
@@ 3.3 Manifold-Informed Regularization
This penalizes minimal-drift updates only when similarity exceeds $\delta$ (default $\delta=0.2$).
@@ 3.4 Overall Process
This creates a lightweight control interface without parameter updates. We inject at every layer using identical interfaces for candidate evaluation and generation, with fixed random matrices of rank $r$ (default $r=8$) to maintain the training-free setting.
@@ 3.5 Vanishing Acceptance of Hard Geometric Constraints
Crucially, $\alpha$ collapses rapidly in high-dimensional latent spaces due to concentration of measure; feasible regions often occupy an exponentially small fraction of the local neighborhood as dimensionality increases (Betancourt, 2017; Goyal and Shetty, 2019; Yallup et al., 2025).
@@ 4.1 Experimental Setup
We use Qwen3 (Yang et al., 2025) as the primary testbed and report main results for the 8B variant; additional scales, 1.7B, 4B and 14B, are deferred to Appendix. To study the interaction with training stages, we also evaluate Phi-4-reasoning (SFT Pre-Training) and Phi-4-reasoning-plus (with RL Post-Training) (Abdin et al., 2025).
@@
We report Pass@$k$ for $k\in\mathcal{K}$ and the normalized area under the Pass@$k$ curve, namely AUC as robustness metrics, along with Avg. Tokens, defined as the average end-to-end tokens per problem required to produce $k$ final trajectories.
@@
where $\mathcal{K}=\{k_{0}<\dots<k_{M}\}$. For code benchmarks, Pass@$k$ is computed using the unbiased estimator of Chen (2021). For math benchmarks, Pass@$k$ is the empirical success rate that at least one of $k$ samples is correct. We use temperature $0.6$ and the default TGR sampling with $K=8$ candidate anchors per chunk boundary, scored by rapid rollouts ($s=32$ on math tasks and $s=64$ on code tasks). We set an upper bound of $L=24$ chunks with maximum chunk length $S=512$, while actual generation terminates early by standard stopping criteria; all reported costs are based on realized token consumption.
@@ 4.4 Interpreting TGR Representations
We find that robustness exhibits a consistent geometric signature, with AUC peaking at an intermediate mean geodesic curvature $\kappa$ that is neither overly rigid nor overly erratic and increasing with the effective candidate-set size proxy $N_{\mathrm{eff}}$. This pattern supports the claim that TGR’s advantage stems from maintaining genuinely distinct search paths rather than redundant variations.